OpenStudy (anonymous):

I'm trying to solve this using the elimination process: 3x+(5/2)y=21 x+2y=14 have i done it right so far? I'm new to the elimination process.. 3x+(5/2)y=21 -(x+2y=14) 2x + 1/2y = 7

OpenStudy (texaschic101):

no... lets get rid of the fraction in the first equation by multiplying it by 2.. (2)3x + (2)5/2y = (2)21 6x + 5y = 42(thats better) Now the problem reads.... 6x + 5y = 42 x + 2y = 14 --->(-6)x + 2y = 14 ------------- 6x + 5y = 42 -6x - 12y = - 84 (result of multiplying by -6) --------------add 0 - 7y = - 42 -7y = - 42 (divide by -7 to get y by itself) -7y/-7 = -42/-7 y = 6 now sub 6 in for y in either of the original equations... x + 2y = 14 x + 2(6) = 14 x + 12 = 14 (subtract 12 from both sides to get x by itself) x + 12 - 12 = 14 - 12 x = 2 now check by subbing in your known variables... x + 2y = 14 2 + 2(6) = 14 2 + 12 = 14 14 = 14 (correct) What you have to try to do in elimination is eliminate " get rid of " one of the variables, so you can find the other variable. Sometimes you have to do this by multiplying the equation by a number so that the variables will cancel each other out. That is why I multiplied the second equation by a -6. And as you can see, it made the x's cancel out and I was able to solve for y. Once finding y, I then substituted what I got for y into the second equation and found x. You can substitute into either one of the equations and you will get the answer. Then I checked my answer by subbing in the known variables and if they equal, which they did, then the problem is correct.

OpenStudy (anonymous):

Thank you ! (:

OpenStudy (texaschic101):

your quite welcome....thats what I am here for